A(0,9), B(8,15) and C(5,k). Find k if A, B and C are collinear points.
Answer by Alan3354(69443)

A(0,9), B(8,15) and C(5,k). Find k if A, B and C are collinear points.
--------------
The x from A to B is 8, from A to C is 5.
--> 5/8 of the distance from A
------
y from A to B is 6
(5/8)*6 = 15/4
k = 9 + 15/4 = 12.75